2(x-1)/3 + 3x/4 = 0

Answer by stanbon(75887)

2(x-1)/3 + 3x/4 = 0
------------
[2x-2)]/3 + (3x)/4 = 0
-----
Multiply thru by 12 to get:
8x-8 + 9x = 0
17x = 8
x = 8/17
---------
Cheers,
Stan H.
-----------